# Mamba-FSCIL: Dynamic Adaptation with Selective State Space Model for Few-Shot Class-Incremental Learning

### v1

@@ Abstract
Few-shot class-incremental learning (FSCIL) confronts the challenge of integrating new classes into a model with minimal training samples while preserving the knowledge of previously learned classes. Traditional methods widely adopt static adaptation relying on a fixed parameter space to learn from data that arrive sequentially, prone to overfitting to the current session. Existing dynamic strategies require the expansion of the parameter space continually, leading to increased complexity. To address these challenges, we integrate the recently proposed selective state space model (SSM) into FSCIL. Concretely, we propose a dual selective SSM projector that dynamically adjusts the projection parameters based on the intermediate features for dynamic adaptation. The dual design enables the model to maintain the robust features of base classes, while adaptively learning distinctive feature shifts for novel classes. Additionally, we develop a class-sensitive selective scan mechanism to guide dynamic adaptation. It minimizes the disruption to base-class representations caused by training on novel data, and meanwhile, forces the selective scan to perform in distinct patterns between base and novel classes. Experiments on miniImageNet, CUB-200, and CIFAR-100 demonstrate that our framework outperforms the existing state-of-the-art methods. The code is available at https://github.com/xiaojieli0903/Mamba-FSCIL.
@@ 1 Introduction
Figure 1: A sketch of our motivation. (a) Static adaptation with MLP is inadequate to separate the features belonging to different sessions that arrive incrementally, prone to overfitting to the current session classes; (b) Our method adopts selective SSM such that the projection learns an adaptive feature shift dependent on the input. Different color refers to data from different sessions.
@@
Deep neural networks often struggle with adapting to new scenarios from limited data without forgetting previous knowledge. This is particularly challenging in real-time environments like robotics and autonomous driving, where data arrives incrementally and new categories often have few samples to learn. To address this challenge, the concept of few-shot class-incremental learning (FSCIL) was introduced [57, 76, 80]. In FSCIL, a model is trained on a base session with adequate labels and samples. Subsequently, it must adapt to new classes encountered in later sessions, each featuring only a few examples, without access to the data from previous sessions. Critically, the model needs to accurately classify images from all encountered classes in each session’s evaluation, balancing the adaptation to new classes with retaining old knowledge.
@@
Various methods have been developed to address FSCIL. Data-based methods [29, 1, 42] replay data subsets or create synthetic samples to reinforce memory and mitigate catastrophic forgetting. Optimization-based techniques [7, 80, 56, 23, 33, 22, 5, 3, 10, 74, 9, 81] utilize advanced loss functions, regularization, and pseudo-feature generation to separate old-class and new-class features. However, most of these methods rely on static adaptation, which uses a pre-defined network architecture that does not change dynamically in response to new information. For example, as shown in Figure 1-(a), multi-layer perceptrons (MLPs) are usually used to project features from the frozen backbone to prototype classifiers [22, 42, 70]. While computationally efficient, MLP cannot dynamically adjust its parameters as class distributions evolve over different learning sessions. As a result, it has insufficient capacity to manage the delicate balance between preserving old knowledge (stability) and accommodating new ability (plasticity), leading to inferior performance on new classes or catastrophic forgetting of old ones. Dynamic network-based methods [57, 67, 66, 68, 76] offer an alternative by expanding architectures with each incremental session. They introduce extra task-specific parameters while preserving the parameters related to the old classes. While enhancing the capacity to incorporate new knowledge, these methods often require continuous structural changes and parameter expansion, leading to increased complexity and resource demands in incremental training.
@@
Recently, Selective State Space Models (SSMs) [17], as known as the Mamba model, have emerged as a compelling alternative to the self-attention mechanisms [58, 11, 32] for efficiently handling long sequences. Mamba introduces a selective scan mechanism that dynamically computes its system matrices based on the input context, allowing for the effective handling of diverse and evolving data characteristics. This dynamic adaptation enables Mamba to adjust its operation parameters according to the needs of different input distributions, increasing the compatibility between old and new classes. Moreover, the 2D Selective Scan (SS2D) proposed for vision tasks [31, 55, 6, 47, 30] further enhances Mamba by spatially arranging image patches in four directions to ensure that each element of the feature map integrates information from all the other locations in different directions, thus creating a global receptive field without increasing linear computational complexity.
@@
Inspired by the appealing merits of SSMs, this study explores how to fulfill the potential of Mamba to facilitate FSCIL. Unlike traditional FSCIL methods that rely on static structures or continually expanding parameters, Mamba is able to learn adaptive representation across old and new classes with a consistent parameter space dependent on the input. As shown in Figure 1-(b), by employing a Selective SSM-based projection layer that dynamically adjusts its parameters based on the input characteristics, the model can maintain the discriminative features of base classes, while compatibly integrating new classes without the need of expanding the parameter space. This not only preserves model efficiency but also minimizes the risk of overfitting, as the model does not accumulate excessive parameters that specialize narrowly on the training data of a particular session. Additionally, Vision Mamba’s [31, 83, 25, 65, 37] proficiency in processing visual tokens can capture fine-grained details of images, which is particularly beneficial when novel classes only have limited training samples, thereby enhancing the model’s perception of incremental classes through selective focus on the relevant image patches.
@@
  • Our approach is thoroughly evaluated on standard FSCIL benchmarks including CIFAR-100, CUB-200, and miniImageNet. Our results achieve state-of-the-art performances, demonstrating effectiveness in improving and preserving base-class accuracy and enhancing the adaptability to novel classes.
2 Related Works
2.1 Few-shot Class-Incremental Learning
@@
FSCIL combines aspects of few-shot learning [45, 59], which focuses on learning from a small number of samples, with class-incremental learning [4, 26, 46], which involves adding new classes to a model’s knowledge base over time without retraining from scratch. FSCIL faces the key challenges: (1) Catastrophic Forgetting. Learning new classes without accessing the data of old classes can cause the model to forget previously acquired knowledge  [38, 15, 72]. (2) Data Scarcity. Limited data availability hampers the model’s ability to learn robust and generalizable representations and increases the overfitting risk [51, 54]. (3) Stability-Plasticity Dilemma. The model must seek a delicate balance, maintaining the knowledge of old classes while adapting to the new ones.
@@
Various methods have been proposed to address the challenges. Data-based strategies mitigate catastrophic forgetting by replaying a subset of real data or generating synthetic data [29, 1, 42]. Optimization-based methods adopt techniques like meta-learning [74, 9, 81], pseudo-feature generation [8, 80], advanced loss functions or regularizers [56, 23, 33, 5, 3, 52, 2, 84, 16], well-designed classifiers [22, 70, 71, 82, 62], and knowledge distillation [57, 10, 7, 78]. For example, C-FSCIL [22] employs a mapping strategy that transforms input images into quasi-orthogonal prototypes to enhance class separability. NC-FSCIL [70] introduces a fixed equiangular tight frame classifier and trains a projection layer to attract the backbone features to these ideal prototypes, thus mitigating inter-class interference. However, most of these methods are based on static adaptation, whose parameters are not aware of the input, increasing the difficulty of striving for a comprehensive ability in both old and new classes. Existing dynamic adaptation methods dynamically adjust the model structure or interrelationships between prototypes [57, 67, 66, 68, 76]. TOPIC [57] uses a neural gas network to integrate new nodes and edges for new classes. CEC [76] employs graph models to evolve the classifier’s topology as new classes emerge. Dynamic networks such as LEC-Net [67] DER [66], and DSN [68] introduce task-specific parameters for new tasks while preserving those related to the old tasks.
@@ 2.2 State Space Models
Transformers [32, 11, 21, 12], despite their proficiency in handling long-range dependencies, are hindered by their quadratic computational cost with respect to input size, particularly for long sequences or high-resolution images. In contrast, State Space Models (SSMs) provide an efficient alternative, capable of processing extensive sequences with significantly reduced computational demands [18, 19, 53, 44, 50, 41, 13, 61, 39, 63, 34]. SSMs involve recurrent updates over a sequence through hidden states, maintaining context in hidden states, and updating outputs by integrating these states with incoming inputs. Structured state space models (S4) [18, 19] exemplify this efficiency with a parameterized linear recurrent neural network. Subsequent research further enhances computational efficiency and expressive capacity of SSMs across various applications [19, 53, 44, 50, 41].
The recently proposed Mamba [17] enhances S4 with a data-dependent selective mechanism to selectively propagate or forget information according to the sequential input tokens and proposes a hardware-aware algorithm to boost efficiency. Mamba has also been adopted to process non-sequential input such as images and videos in vision tasks  [31, 83, 47, 30, 65, 37, 25]. For example, VMamba [31] introduces a cross-scan mechanism that processes images in both horizontal and vertical dimensions, allowing feature map elements to integrate information from all directions and providing a global receptive field without increasing linear computational complexity.
Although the Mamba model has been extensively applied in various domains such as text [17, 13, 18], vision [83, 31, 25, 35], audio [14], its potential to mitigate model forgetting has not yet been explored. This study is the first attempt to apply Mamba in FSCIL, effectively harnessing its capabilities of dynamic adaptation and sequence modeling to achieve state-of-the-art performance in FSCIL.
@@ 3.2 Mamba
Selective State Space Models (S6)
are the core component of Mamba, enhancing S4 by introducing a selective and dynamic mechanism for the interactions between sequential states [17]. Unlike conventional SSMs that utilize constant operation matrices that are agnostic of input features, S6 models utilize input-dependent parameters that calculate parameters ${\mathbf{B}}$, ${\mathbf{C}}$, and $\mathbf{\Delta}$ directly from the input sequence ${\bm{x}}$ through linear projections, enabling a dynamic and context-aware adaptation to the unique characteristics of the data.
2D Selective Scan (SS2D)
enhances Mamba’s ability to effectively capture spatial information for vision tasks [31]. SS2D arranges the feature patches in four different directions (top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right), creating four separate sequences. Each sequence is processed by the S6 module before merging to reconstruct the 2D feature map. Given the input feature ${\bm{z}}$, the output feature $\overline{{\bm{z}}}$ of SS2D is formulated as:
$\overline{{\bm{z}}}=\text{SS2D}({\bm{z}})=\sum_{i=1}^{4}\text{S6}(\text{scan}({\bm{z}},i)),$ (5)
where $\text{scan}({\bm{z}},i)$ converts feature patches into sequence according to direction $i$. This approach ensures the comprehensive integration of contextual information from all directions, enhancing the model’s global receptive field.
@@ 4 Mamba-FSCIL
Figure 2: Mamba-FSCIL framework. The macro architecture (a) is composed of a backbone network, a dual selective SSM projector consisting of an identity branch and two selective SSM branches with the identical structure of (b), and a classifier. Our class-sensitive selective scan (c) performs $\mathcal{L}_{\text{supp}}$ and $\mathcal{L}_{\text{sep}}$ in the selective SSM branch to guide the dynamic adaptation in incremental session’s training.
@@ 4.1 Macro architecture
The architecture of Mamba-FSCIL is composed of a backbone network $f$, a dual selective SSM projector, and a linear classifier. The backbone network $f$ processes the input ${\bm{x}}$ to generate intermediate features ${\mathbf{F}}=f({\bm{x}})\in\mathbb{R}^{N\times D\times H\times W}$, where $N$, $D$, $H$, and $W$ denote the batch size, number of channels, height, and width, respectively. It is only trainable in the base session to learn a powerful feature extractor using the extensive training data of the base session. We design a dual selective SSM projector that maps the frozen intermediate features into the final representation for dynamic adaptation. The classifier adopts a fixed simplex equiangular tight frame (ETF) following [70, 79, 27, 49, 64, 36]. It provides an ideal feature space arrangement with the maximal equiangular separation among classes [69, 40].

### v2

@@ Abstract
Few-shot class-incremental learning (FSCIL) confronts the challenge of integrating new classes into a model with minimal training samples while preserving the knowledge of previously learned classes. Traditional methods widely adopt static adaptation relying on a fixed parameter space to learn from data that arrive sequentially, prone to overfitting to the current session. Existing dynamic strategies require the expansion of the parameter space continually, leading to increased complexity. In this study, we explore the potential of Selective State Space Models (SSMs) for FSCIL, leveraging its dynamic weights and strong ability in sequence modeling to address these challenges. Concretely, we propose a dual selective SSM projector that dynamically adjusts the projection parameters based on the intermediate features for dynamic adaptation. The dual design enables the model to maintain the robust features of base classes, while adaptively learning distinctive feature shifts for novel classes. Additionally, we develop a class-sensitive selective scan mechanism to guide dynamic adaptation. It minimizes the disruption to base-class representations caused by training on novel data, and meanwhile, forces the selective scan to perform in distinct patterns between base and novel classes. Experiments on miniImageNet, CUB-200, and CIFAR-100 demonstrate that our framework outperforms the existing state-of-the-art methods. The code is available at https://github.com/xiaojieli0903/Mamba-FSCIL.
Index Terms:
Few-Shot Class Incremental Learning, Selective State Space Model, Dynamic Adaptation, Class Incremental Learning
@@ 1 Introduction
Deep neural networks often struggle with adapting to new scenarios from limited data without forgetting previous knowledge. This is particularly challenging in real-time environments like robotics and autonomous driving, where data arrives incrementally and new categories often have few samples to learn. To address this challenge, the concept of few-shot class-incremental learning (FSCIL) was introduced [1, 2, 3]. In FSCIL, a model is trained on a base session with sufficient labels and samples. Subsequently, it must adapt to new classes encountered in later sessions, each containing only a few examples, without access to the data from previous sessions. Critically, the model needs to accurately classify images from all encountered classes in each session’s evaluation, balancing the adaptation to new classes with retaining knowledge of previously learned classes.
@@
Various methods have been developed to address FSCIL [4, 5]. Data-based methods [6, 7, 8] replay data subsets or create synthetic samples to reinforce memory and mitigate catastrophic forgetting. Optimization-based techniques [9, 10, 11, 12, 13, 14, 15, 16, 17, 18] utilize advanced loss functions, regularization, and pseudo-feature generation to separate old-class and new-class features. However, most of these methods rely on static adaptation, which uses a pre-defined network architecture that does not change dynamically in response to new information. For example, as shown in Figure 1 (a), multi-layer perceptrons (MLPs) are usually used to project features from the frozen backbone to prototype classifiers [8, 16, 19]. While computationally efficient, MLP cannot dynamically adjust its parameters as class distributions evolve over different learning sessions. As a result, it has insufficient capacity to manage the delicate balance between preserving old knowledge (stability) and accommodating new ability (plasticity), leading to inferior performance on new classes or catastrophic forgetting of old ones. Dynamic network-based methods [1, 2, 20, 21, 22, 23, 24, 25, 26] offer an alternative by expanding architectures with each incremental session. They introduce extra task-specific parameters while preserving the parameters related to the old classes. While enhancing the capacity to incorporate new knowledge, these methods often require continuous structural changes and parameter expansion, leading to increased complexity and resource demands in incremental training.
Figure 1: A sketch of our motivation. (a) Static Adaptation with MLP: MLPs apply the same transformation to all inputs, which is inadequate for separating evolving features from different sessions and prone to overfitting to current session classes; (b) Dynamic Adaptation with Selective SSMs: Selective SSMs dynamically adjust parameters based on input $x_{t}$, generating input-dependent parameters ($\bm{\Delta}_{t}$, ${\mathbf{B}}_{t}$, ${\mathbf{C}}_{t}$) to learn adaptive feature shifts in the output $y_{t}$. Different colors (red, green, purple) represent data from different sessions.
@@
Recently, Selective State Space Models (SSMs) [27], as known as the Mamba model, have emerged as a compelling alternative to self-attention mechanisms [28, 29, 30] to efficiently handle long sequences. Mamba introduces a selective scan mechanism that dynamically computes its system matrices based on the input context, allowing for the effective handling of diverse and evolving data characteristics. This dynamic adaptation enables Mamba to adjust its operation parameters according to the needs of different input distributions, increasing the compatibility between old and new classes. Moreover, the 2D Selective Scan (SS2D) proposed for vision tasks [31, 32, 33, 34] further enhances Mamba by spatially arranging image patches in four directions to ensure that each element of the feature map integrates information from all other locations in different directions, thus creating a global receptive field without increasing linear computational complexity.
@@
Inspired by the appealing merits of SSMs, this study explores how to fulfill the potential of Mamba to facilitate FSCIL. Unlike traditional FSCIL methods that rely on static structures or continually expanding parameters, Mamba is able to learn adaptive representation across old and new classes with a consistent parameter space dependent on the input. As shown in Figure 1 (b), by employing a Selective SSM-based projection layer that dynamically adjusts its parameters based on input characteristics, the model can maintain the discriminative features of the base classes, while compatibly integrating new classes without the need of expanding the parameter space. This not only preserves model efficiency but also minimizes the risk of overfitting, as the model does not accumulate excessive parameters that specialize narrowly on the training data of a particular session. Additionally, Vision Mamba’s [31, 35, 36, 37, 38, 39] proficiency in processing visual tokens can capture fine-grained details of images, which is particularly beneficial when novel classes only have limited training samples, thereby enhancing the model’s perception of incremental classes through selective focus on the relevant image patches.
@@
  • Our approach is thoroughly evaluated on standard FSCIL benchmarks, including CIFAR-100, CUB-200, and miniImageNet. Our results achieve state-of-the-art performances, demonstrating effectiveness in improving and preserving base-class accuracy and enhancing the adaptability to novel classes.
2 Related Work
2.1 Few-Shot Learning (FSL)
FSL enables models to classify new classes with limited training samples [40, 41, 42, 43], and can be divided into four main strategies: data augmentation, metric-based, model-based, and optimization-based approaches [44]. Data augmentation methods reduce overfitting and enhance generalization by increasing data diversity through transforming existing samples or synthetic data generation using models like VAEs or GANs [45]. Metric-based methods classify objects by computing similarity or distance between samples in a learned embedding space, with common approaches including Siamese Networks [46], Matching Networks [47], and Prototypical Networks [48]. Model-based methods design specific architectures for FSL challenges, such as Memory-Augmented Neural Networks[49] and Meta Networks [50], to store crucial class information and learn generalizable meta-knowledge. Optimization-based methods aim for fast model adaptation with few-shot data. Model-Agnostic Meta-Learning (MAML) [51] is a common optimization algorithm that trains the model’s initial parameters using various datasets to ensure a peak performance when tackling new tasks. Reptile [52] extends MAML by simplifying gradient calculations, enhancing computational efficiency.
While FSL focuses on generalizing to new classes with limited data, it does not prioritize retaining knowledge of base classes. In contrast, FSCIL requires models to continuously learn new classes with limited samples while preserving the knowledge of previously learned classes.
2.2 Class-Incremental Learning (CIL)
CIL enables models to learn new classes over time without forgetting previously acquired knowledge [53, 54, 55, 56, 57, 9, 24, 58]. CIL approaches include non-exemplar-based and exemplar-based methods. Non-exemplar-based methods focus on preserving existing knowledge without relying on stored examples. Methods like EWC [59], MAS [60], and SI [61] measure the importance of parameters and apply regularization to prevent significant changes in critical parameters. Additionally, some methods dynamically change the model architecture, using network expansion and pruning to accommodate new tasks [62, 63, 22]. Exemplar-based methods, e.g., iCaRL [55], combine exemplar replay with knowledge distillation, while BiC [64] forms a validation set from saved exemplars and adjusts outputs via a scaling layer. GEM [65, 66] uses exemplars for gradient projection to prevent knowledge overwriting. Additional approaches include storing embeddings rather than raw images [67], using generative models for replay [68], and implementing task-wise adaptation and output normalization [69, 70, 71, 72].
While CIL typically has sufficient training data in incremental sessions, FSCIL complicates this by limiting training samples, making the prevention of catastrophic forgetting even more difficult.
2.3 Few-shot Class-Incremental Learning (FSCIL)
@@
FSCIL [4, 5] combines the aspects of FSL and CIL, continuously learning new classes with a small number of samples to a model’s knowledge base over time without retraining from scratch. FSCIL faces the key challenges: (1) Catastrophic Forgetting. Learning new classes without accessing the data of old classes can cause the model to forget previously acquired knowledge  [73, 74, 75]. (2) Data Scarcity. Limited data availability hampers the model’s ability to learn robust and generalizable representations and increases the overfitting risk [48, 76]. (3) Stability-Plasticity Dilemma. The model must seek a delicate balance, maintaining the knowledge of old classes while adapting to the new ones [77, 78, 79].
@@
Various methods have been proposed to address the challenges. Data-based strategies mitigate catastrophic forgetting by replaying a subset of real data or generating synthetic data [6, 7, 8]. Optimization-based methods adopt techniques like meta-learning [20, 17, 18], pseudo-feature generation [80, 3], advanced loss functions or regularizers [9, 14, 15, 11, 12, 81, 82, 83, 84], well-designed classifiers [16, 19, 85, 86, 87], and knowledge distillation [1, 13, 10, 88]. For example, C-FSCIL [16] employs a mapping strategy that transforms input images into quasi-orthogonal prototypes to enhance class separability. NC-FSCIL [19] introduces a fixed equiangular tight frame classifier and trains a projection layer to attract the backbone features to these ideal prototypes, thus mitigating inter-class interference. However, most of these methods are based on static adaptation, whose parameters are not aware of the input, increasing the difficulty of striving for a comprehensive ability in both old and new classes. Dynamic adaptation methods dynamically adjust the model structure or interrelationships between prototypes [1, 21, 22, 23, 2, 89]. TOPIC [1] uses a neural gas network to integrate new nodes and edges for new classes. CEC [2] employs graph models to evolve the classifier’s topology as new classes emerge. Dynamic networks such as LEC-Net [21], DER [22], DSN [23] and FeSSS [89] introduce task-specific parameters for new tasks while preserving those related to old tasks.
@@ 2.4 State Space Models
Transformers [28, 29, 30], despite their proficiency in handling long-range dependencies, are hindered by their quadratic computational cost with respect to input size and large inference-time memory requirements from the key-value cache, particularly for long sequences or high-resolution images. In contrast, State Space Models (SSMs) provide an efficient alternative, capable of processing extensive sequences with significantly reduced computational demands [90, 91, 92, 93, 94, 95, 96, 97, 98, 99]. SSMs involve recurrent updates over a sequence through hidden states, maintaining context in hidden states, and updating outputs by integrating these states with incoming inputs. Structured state space models (S4) [90, 91] exemplify this efficiency with a parameterized linear recurrent neural network. Subsequent research further enhances computational efficiency and expressive capacity of SSMs across various applications [91, 92, 93, 94, 95].
Mamba [27, 100] enhances S4 with a data-dependent selective mechanism to selectively propagate or forget information according to sequential input tokens and proposes hardware-aware algorithms to boost efficiency. Mamba has also been adopted to process non-sequential input such as images and videos in vision tasks  [31, 35, 33, 34, 37, 38, 36, 39]. Among them, VMamba [31] introduces a cross-scan mechanism that processes images in both horizontal and vertical dimensions, allowing feature map elements to integrate information from all directions and providing a global receptive field without increasing linear computational complexity.
Although the Mamba model has been extensively applied in various domains such as text [27, 96, 90, 101], vision [35, 31, 36, 102], audio [103], its potential to mitigate model forgetting has not yet been explored. This study is the first attempt to apply Mamba in FSCIL, effectively harnessing its capabilities of dynamic adaptation and sequence modeling to achieve state-of-the-art performance.
@@ 3.2 Mamba
Selective State Space Models (S6) are the core component of Mamba, enhancing S4 by introducing a selective and dynamic mechanism for the interactions between sequential states [27]. Unlike conventional SSMs that utilize constant operation matrices that are agnostic of input features, S6 models utilize input-dependent parameters that calculate parameters ${\mathbf{B}}$, ${\mathbf{C}}$, and $\mathbf{\Delta}$ directly from the input sequence ${\bm{x}}$ through linear projections, enabling a dynamic and context-aware adaptation to the unique characteristics of the data.
2D Selective Scan (SS2D) enhances Mamba’s ability to effectively capture spatial information for vision tasks [31]. SS2D arranges the feature patches in four different directions (top-left to bottom-right, bottom-right to top-left, top-right to bottom-left, and bottom-left to top-right), creating four separate sequences. Each sequence is processed by the S6 module before merging to reconstruct the 2D feature map. Given the input feature ${\bm{z}}$, the output feature $\overline{{\bm{z}}}$ of SS2D is formulated as:
$\overline{{\bm{z}}}=\text{SS2D}({\bm{z}})=\sum_{i=1}^{K}\text{S6}(\text{scan}({\bm{z}},i)),$ (5)
where $\text{scan}({\bm{z}},i)$ converts feature patches into a sequence according to direction $i$, and $K$ denotes the number of scan paths (default is 4). This approach ensures the comprehensive integration of contextual information from all directions, enhancing the model’s global receptive field.
@@ 4 Mamba-FSCIL
Figure 2: Mamba-FSCIL framework. The macro architecture (a) is composed of a backbone network $f$, a dual selective SSM projector consisting of an identity branch $p$ and two selective SSM branches ($g^{\text{base}}$ for the base session and $g^{\text{inc}}$ for incremental sessions) with the identical structure of (b), and a classifier. Our class-sensitive selective scan (c) performs $\mathcal{L}_{\text{supp}}$ and $\mathcal{L}_{\text{sep}}$ in the incremental selective SSM branch $g^{\text{inc}}$ to guide the dynamic adaptation in incremental session’s training.
@@ 4.1 Macro Architecture
The architecture of Mamba-FSCIL, illustrated in Fig. 2 (a), is composed of a backbone network $f$, a dual selective SSM projector, and a linear classifier. The backbone network $f$ processes the input ${\bm{x}}$ to generate intermediate features ${\mathbf{F}}=f({\bm{x}})\in\mathbb{R}^{N\times D\times H\times W}$, where $N$, $D$, $H$, and $W$ denote the batch size, number of channels, height, and width, respectively. It is only trainable in the base session to learn a powerful feature extractor from the extensive base session data and is then frozen during incremental sessions, a common practice in FSCIL [2, 19, 104, 83] to prevent catastrophic forgetting caused by fine-tuning on novel classes.
We develop a dual selective SSM projector that maps the frozen intermediate features into the final representation for dynamic adaptation. It consists of three branches: a base session branch $g^{\text{base}}$, an incremental session branch $g^{\text{inc}}$, and an identity branch $p$. The base session branch and incremental session branch have the same structure shown in Fig. 2 (b). The base session branch is responsible for projecting the features of the base classes into a stable and robust representation, which is frozen during the incremental sessions. The incremental session branch dynamically adapts to new classes by learning feature shifts while preserving the previously learned knowledge. The details of this projector will be introduced in Section 4.2.
To further enhance dynamic adaptation, a class-sensitive selective scan mechanism is proposed. This mechanism guides the incremental session branch $g^{\text{inc}}$ in adapting to new classes with minimal interference to the representations of base classes. As shown in Fig. 2 (c), it employs suppression and separation losses to manage the impact of new classes and enforce distinct patterns of feature adaptation between base and novel classes. This mechanism will be detailed in Section 4.3.
The classifier adopts a fixed simplex equiangular tight frame (ETF) following [19, 105, 106, 107, 108, 109]. It provides an ideal feature space arrangement with the maximal equiangular separation among classes [110, 111]. In Section 4.4, we will introduce the training and optimization process across sessions.